### v3

# X-TIME: An in-memory engine for accelerating machine learning on tabular data with CAMs

Abstract
Structured, or tabular, data is the most common format in data science. While deep learning models have proven formidable in learning from unstructured data such as images or speech, they are less accurate than simpler approaches when learning from tabular data. In contrast, modern tree-based Machine Learning (ML) models shine in extracting relevant information from structured data. An essential requirement in data science is to reduce model inference latency in cases where, for example, models are used in a closed loop with simulation to accelerate scientific discovery. However, the hardware acceleration community has mostly focused on deep neural networks and largely ignored other forms of machine learning. Previous work has described the use of an analog content addressable memory (CAM) component for efficiently mapping random forests. In this work, we focus on an overall analog-digital architecture implementing a novel increased precision analog CAM and a programmable network on chip allowing the inference of state-of-the-art tree-based ML models, such as XGBoost, CatBoost, and others. Results evaluated in a single chip at 16nm technology show 119$\times$ higher throughput at 9740$\times$ lower latency compared with a state-of-the-art GPU, with a 19W peak power consumption.
Index Terms:
In-memory computing, tree-based ML, content addressable memories
I Introduction
Extracting relevant information from structured (i.e., tabular) data is paramount in practical data science. Tabular data consist of a matrix of samples organized in rows, each of which has the same set of features in its columns, and is commonly used for medical, financial, scientific, and networking applications, enabling efficient search operations on large-scale datasets. While deep learning has shown impressive improvements in accuracy for unstructured data, such as image recognition or machine translation, tree-based machine learning (ML) models still outperform neural networks in classification and regression tasks on tabular data [1]. This difference in model accuracy is due to robustness to uninformative features of tree-based models, bias towards a smooth solution of neural networks models, and in general the need for rotation invariant models where the data is rotation invariant itself, as in the case of tabular data [1]. With high model accuracy, ease of use, and explainability [2] tree-based ML models are greatly preferred by the data science community [3], with $>74\%$ of data scientists preferring to use these models while $<40\%$ choose neural networks in a recent survey.
However, tree-based models have garnered little attention in the architecture and custom accelerator research fields, particularly compared to deep learning. While small models are easily executed, performance issues for larger tree-based ML models arise due to limited inference latency both on CPU and GPUs [4, 5] originating from irregular memory accesses and thread synchronization issues. As large tree-based models are finding performance limits in existing hardware, larger models ($>$1M nodes) are demonstrating increasingly compelling uses in practical environments where high throughput and moderate latency are desired. For example, the winner of a recent IEEE-CIS Fraud detection competition used a tree-based ML model with 20M nodes [6]. Moreover, reducing inference latency is critical in scientific applications where for use in real-time processing, model latency must be $\sim$100 ns [7]. Similar low latency requirements for model inference arise in a real-time in-the-loop decision, data filtering systems, or streaming data scenarios, which are common across financial, medical, and cybersecurity domains [8, 6] where tabular data and tree-based ML also dominate. Thus, these models are reaching sizes and complexity levels for meaningful real-world tasks, as they are hitting limits on inference latency and throughput in current systems. Recently, custom accelerators for tree-based ML inference have been developed based either on FPGA [7, 8] or custom ASICs [9, 5]. A common architecture in most of the accelerators mentioned above implements look-up tables (LUTs) storing the attributes of a decision tree (DT) accessed multiple times during inference, with reads based on decision paths taken during the tree traversal. Along with low latency inference, an efficient hardware representation of trees is necessary to implement the required very large models.
One promising new architectural approach for improving tree-based model inference is in-memory computing, where data are stored and processed in the same location. This approach eliminates access, transfer, and processing of data in a separate compute unit [10], reducing latency. Previous research has shown that these devices can efficiently perform multiplication functions in crossbar architectures [11]. However, crossbar architectures are limited to matrix operations, which are not useful for inference on tree-based models. Another class of in-memory computing primitives is Content Addressable Memories (CAMs) [12, 13], high-speed and highly parallel circuits that search for a given input in a table of stored words and return the matched search input’s location. More recently, analog CAMs have been proposed [14] which leverage the analog states in memristor devices. An analog CAM cell performs in situ comparisons of stored (analog) ranges with (analog) inputs, by computing if the input is within the range. Effectively, analog CAM arrays perform a massively parallel comparison between an input vector of analog values and stored words of analog ranges. Given that DTs are a set of conditional branches, analog CAMs naturally map DTs and perform highly parallel inference of decision forests [15]. Previous work showed the potential of this approach but was limited to component-level estimation, 4-bit precision, and random forest models [15].
This work presents X-TIME, an in-memory engine for accelerating tree-based ML models with analog CAMs. In this work we demonstrate the need for a holistic architecture design, 8-bit operation, and support for multiple tree-based ML models, overcoming the limitation of previous work [15]. The major contributions of this paper compared to previous work [15] are (1) a novel analog CAM cell design and operation to perform decision tree inference at doubled (i.e. 8-bit) memristor precision (i.e. 4-bit), (2) an in-core and in-network leave accumulation architecture to seamlessly perform inference of a large number of trees in parallel, (3) a programmable Network-on-Chip (NoC) to operate any tree-based ML model regardless of the ensemble reduction operation, (4) a compiler supporting the mapping of multiple tree-based ML models and (5) a rigorous characterization method and scaling comparison with respect to GPUs, FPGAs, and conventional digital accelerators. Our simulated evaluation shows up to 119$\times$ increased throughput at 9740$\times$ decreased latency in a single 19W chip.
II Background
Figure 1: Schematic illustration of (a) a decision tree and (b) a decision tree ensemble.
II-A Tree-based Machine Learning
A DT is a regression and classification ML model consisting of a set of nodes arranged in a tree-like structure. Each node can have multiple children nodes, but the most common case is a binary tree, where each node has only two children. The first node is usually called the root. Each node $j$ consists of a conditional branch where a single input feature element $f_{i}$ is compared with a trainable threshold $T_{j}$. The exception to this is the last node, or leaf, which stores the model’s prediction for a given input, and which can be a class (classification) or a scalar value (regression). A balanced tree can be defined by its depth $D$, which is equal to the maximum number of nodes in a root-to-leaf traversal. Fig. 1(a) shows an example of a DT with $L=4$ leaves and $D=2$.
Multiple DTs can be trained together as an ensemble for increased accuracy in classification and regression tasks. Fig. 1(b) shows a conceptual representation of a DT ensemble, where multiple trees are executed in parallel and their output combined with a reduction operation yields the model’s final result. Examples of DT ensembles model include Random Forests (RF) [16], eXtreme Gradient Boosting (XGBoost) [17], and Categorical Boosting (CatBoost) [18]. In the case of RF, the reduction operation consists of a majority voting, while in all the others leaf values of each DT are summed.
II-B Tree-based ML model inference on GPU
While training DT models is relatively fast when compared to, for example, a large deep neural network, the computations involved in inference are very irregular and not particularly suitable for CPUs and GPUs, becoming inefficient for very large models, with $>1$ M nodes [6]. While CPUs can run irregular control paths fairly well, their parallelism is limited. GPUs have a very large parallelism, but suffer from the irregularity of memory access patterns, as will be explained in detail below. Recent NVIDIA libraries implementations have shown order of magnitude improvements compared to CPU and previous GPU implementations  [19]. However, three different factors still limit GPU throughput and latency [4], in particular for tree ensembles. First, a high number of uncoalesced memory accesses, due to the low ratio between requested and fetched data, causes low memory efficiency. Trees are usually stored in memory with nodes from different trees interleaved in breadth-first order, resulting in coalesced memory accesses close to the root, where children nodes are more likely to be part of the same memory transaction of the parent node, but uncoalesced memory accesses further from the root. Thus, with increasing tree depth the proportion of coalesced memory accesses decreases and memory system efficiency also decreases. Second, synchronization between threads and the consequent load imbalance due to waiting for the deepest trees limits latency. A common implementation scheme on GPU assigns each tree in the ensemble to a different thread by round-robin. At the end of each thread block, a reduction operation is then performed. However, a given ensemble may contain short and tall trees, with the latter having substantially larger latency. Thus, each thread block must wait until the slowest (tallest) tree has finished processing, increasing overall latency. Finally, with a larger number of trees $N_{trees}$ in the ensemble, the increasing number of thread blocks and the global reduction performed between thread blocks becomes a significant overhead. Multiple techniques have been proposed to increase the throughput of tree ensemble inference [19, 4], mainly by re-organizing the memory - for example, for sparse trees and scheduling of threads - but with limited impact on latency. In this work, we focus on addressing these limitations with an in-memory computing architecture using content-addressable memories.
Figure 2: Comparison of (a) a conventional CAM/TCAM and (b) an analog CAM
II-C Analog Content Addressable Memory
CAMs are specialized memory devices that search a table of stored words (rows) or keys in parallel for a given input query. The CAM search output is a one-hot vector where a high value corresponds to a match between the input data query and stored row word key [12, 13]. In conventional CAMs, this output vector is converted into a single address by a priority encoder. Each CAM cell functions as both memory and bit comparison, and the input query is compared to each CAM word simultaneously to produce a high throughput, low latency few-cycle compare operation at the cost of high power and area. In addition to binary (0 and 1) values, ternary CAMs (TCAMs) have a third wildcard value, X (or “don’t care”), which can be used to compress the stored CAM table or perform partial match searches. When an X is stored or searched, that cell matches irrespective of whether the searched or stored value is 1 or 0. Fig. 2(a) shows a schematic illustration of a conventional CAM/TCAM circuit, where input queries are applied on the column search lines (SLs) and outputs are read along the row ”match lines” (MALs). An XNOR operation occurs between the query values and stored cell values in the corresponding column. The MAL $j_{th}$ row output is produced by a bit-wise AND operation between cells along the row and returns a match (high) or mismatch (low). The red boxes in Fig. 2(a) correspond to mismatches between input and stored value. In practice, MALs are initially pre-charged high and a mismatch in a cell activates a pull-down transistor connected to MAL to discharge it, such that a match is returned only if all pull-down transistors remain off during a search operation.
TCAMs are used ubiquitously in networking applications that require high-throughput and low-latency operations, for example for packet classification and forwarding, and access control lists [12]. While offering high performance, conventional SRAM-TCAMs are power and area-hungry and require 16 transistors. Recently, there has been significant interest in developing TCAMs with memristor devices and resistive non-volatile memories [20, 21, 22] with compact size and low-power operation. As programming these devices is relatively slow, emerging TCAMs have targeted in-memory computing applications [23, 13] where stored value updates are infrequent.
Beyond direct digital memory operation, memristor devices have also demonstrated ranges of stable analog states of $>16$ levels. Leveraging this analog state capability, recent work proposed and demonstrated an analog CAM circuit design and operation [14] (conceptual representation in Fig 2(b)). In an analog CAM, ranges are stored in each cell, an analog input is applied vertically along columns and a match is returned along rows if the inputs are all within the stored range of each cell in a row. An analog CAM row implements
$\text{MAL}_{i}=\bigwedge_{j}{(T_{L,ij}\leq q_{j}<T_{H,ij})}\text{,}$ (1)
where $q$ is the query input vector applied to DL, and $T_{L}$ and $T_{H}$ are the stored range thresholds for the CAM cell range lower and upper bounds, respectively. In the analog CAM circuit, no analog-to-digital conversion is required, eliminating an expensive component which is a limiting bottleneck of other analog in-memory computing primitives [24]. Multiple analog CAM circuits have been realized using different technologies such as memristor devices [14] and ferroelectric memory devices [25]. In such circuits, two analog memory elements represent the lower and upper bound thresholds. Analog CAMs have been shown to map CAM tables more efficiently [14], as well as implement novel functions, such as computing distances for memory-augmented neural networks [25, 22], decision tree inference [15] and pattern learning [26].
Figure 3: DT example from Fig.1(a) with nodes mapped in an analog CAM and leaves mapped into a RAM.
II-D Mapping tree models to analog CAMs
The X-TIME compiler takes as input the description of a DT ensemble in a tabular mode, where each row contains the node attributes (feature ID and value) and the left/right children. The compiler traverses the tree structures extracting all the root-to-leaf paths. Finally, it maps each of them to an ACAM row. As an example, Fig. 3 shows the mapping of Fig. 1(a). Each column corresponds to a different feature of the input vector $q$. Each node is mapped in the column corresponding to the feature used in the node’s comparison. The comparison’s threshold value is programmed in the analog CAM’s lower and/or upper bound, according to the direction, left or right, taken by the comparison’s result for the root-to-leaf path. A “don’t care” is used if a feature is missing, by programming the full range (0 to 1 in our example) in the analog CAM cell. The feature vector (query) is applied on the DL and the output on the MAL is high for the matched branch, i.e. the traversed root-to-leaf path, in a single CAM operation. This mapping requires the number of CAM rows to match the number of tree leaves, and the number of CAM columns to match the number of features $N_{feat}$. MALs can be connected directly to the Word Lines (WLs) of a conventional RAM storing leaf values to return the model’s prediction. Effectively, the analog CAM traverses root-to-leaf paths, and the RAM retrieves leaf values.
Previous results of this approach have shown significant speedups for RF inference [15] and reduced energy consumption compared to both GPU and state-of-art ASICs [27], although with limited architecture investigation for scaling and no flexibility to accommodate different dataset or model types. In this work, we present a complete architecture development and study for handling multiple DT ensemble algorithms such as RF, XGBoost, and CatBoost, and learning tasks such as regression, binary classification, and multi-class classification with the same architecture. In addition, we extend the use of the previously presented analog CAM [14, 15] circuit to support higher precision data by developing a flexible connectivity strategy and bit-arithmetic refactoring. We compare the performance of our architecture on state-of-the-art tabular dataset ML tasks against GPU, FPGA and digital ASIC, showing a large improvement in latency and throughput while using a single chip with limited power consumption.
III Architecture organization
Figure 4: Architecture hierarchy
Fig. 4 shows the internal architecture of an X-TIME single-chip accelerator. The accelerator is composed of a group of cores connected by an H-tree NoC topology converging to a co-processor (CP). Each core contains an analog CAM array acting as a processing element (PE-CAM), buffer, multiple match resolver (MMR), SRAM memory, and accumulator (ACC) (i.e. feedback adder). Typical CAM architectures also contain similar elements but the memory is used as an address encoder rather than for storing leaves [12]. To map models larger than a single array, each PE-CAM has a number of stacked arrays (extended row-wise) and queued arrays (extended column-wise) to handle more and larger words, respectively. Stacked arrays share the same peripherals, namely the Pre-Charger (P-Ch), the Digital Analog Converter (DAC), the Sense Amplifier (SA), and the Match-Line Register (ML-REG). The connection of queued arrays ensures that only previously matched lines are charged by making the ML-REG of array $i$ feed the P-Ch of the array $i+1$.
III-A Inference
X-TIME compiler maps the trained trees to the cores according to Fig. 3. Training is performed using standard libraries such as scikit-learn[28] or XGBoost[17]. XGBoost outputs a table with $N_{leaves}+N_{nodes}$ (where $N_{leaves}$ is the total number of leaves in the ensemble) rows each storing the node ID, left and right child ID, node value, node feature, leaf value (if it is a leaf) and tree ID. Converting this table for CAM usage, we obtain a new table of size $L\times(2N_{feat}+3)$ with each row storing the lower/upper bound for each feature, the leaf value, class ID/label (in the case of multi-class classification) and tree ID, representing the model ensemble for our architecture. The X-TIME compiler assigns trees to cores via round-robin, directly mapping the first $L\times(2N_{feat})$ columns of a tree to the lower and upper bound of the analog CAM circuit, and the third-last column to the SRAM. Class and tree ID are uniquely represented in the core address. The largest ensemble in our architecture is constrained by the maximum number of leaves $N_{leaves,max}$ in all trees and the number of available cores $N_{C}$. If $L<N_{words}/2$, multiple trees can be mapped to a single core, where $N_{words}=N_{stacked}\times H$ is the number of words addressable by each core (with $H$ number of rows in an analog CAM array and $N_{stacked}$ number of arrays stacked).
During inference, a feature vector is sent through the network to a core and searched in the PE-CAM. Each queued analog CAM block receives a different portion of the feature vector which is converted into analog voltage values by a DAC. A shift operation before the conversion is performed to increase operation precision (as discussed in section III-B). A logical AND operation is inherently performed between the queued arrays, similar to MAL segmentation used in modern TCAMs to maintain high throughput where MALs are divided into sections of shorter words with an AND between individual MAL results. Connecting multiple TCAM arrays with common MALs is logically equivalent to a larger TCAM, with the bit-wise AND operation between TCAM cells on the same MAL now extended between arrays. In our case, the analog CAM stores a branch split into multiple analog CAM arrays and reconstructed by means of the logical AND, such that only if all branch nodes arranged in multiple arrays are matched will a match be returned on the corresponding MAL.
The final output of the PE-CAM is stored in a buffer and fed into the MMR which selects only one match signal to proceed logically. The MMR is needed due to the possibility of mapping multiple trees in a single core. In fact, given $N_{trees,core}$ mapped to a core, for each inference, the PE-CAM would return $N_{trees,core}$ matches in the form of a sparse Boolean vector with $N_{trees,core}$ match values (“1”s). The MMR generates $N_{trees,core}$ one-hot Boolean vectors used for directly accessing the SRAM WLs. The SRAM stores the corresponding leaf value of each branch, which is retrieved sequentially and accumulated by the ACC block, which performs a local reduction operation. After $N_{trees,core}$ accumulations are performed, the resulting leaf value is sent as output to the router. Leaf values coming from multiple cores can be summed at the router level as section III-D discusses, performing a shared reduction operation. Finally, the resulting leaf value is propagated to the CP which performs a global reduction and computes the model prediction. The latter is usually a simple operation, such as a majority voting or a threshold comparison.
III-B Analog CAM with improved precision
Figure 5: (a) analog CAM 8-bit macro-cell and (b) analog CAM 4-bit sub-cell circuit.
ML algorithms usually do not require high-precision floating point operations for achieving good results [29]. As will be detailed in section IV we found out that typically 8 bits precision, corresponding to 256 bins for each feature, is required for matching floating point (FP) accuracy inference of tree-based ML models.
While analog CAM circuits offer the possibility of mapping arbitrary ranges, memristor device reliability poses a challenge in supporting more than 16 levels or 4 bits [15]. In the case of crossbar arrays of memristors mapping algebraic matrices, such as a neural network layer, multiple techniques have been proposed for overcoming limited device precision. For example, with bit slicing [11] by using $M$ levels memristor, $N=2^{2M}$ bits can be represented by connecting two devices on the same row. In crossbar arrays, it is also straightforward to arbitrarily adjust the input precision by applying binary pulses and performing SHIFT and ADD operations at the output [11].
Unfortunately, such techniques do not apply to analog CAM arrays. The analog CAM encodes a unary representation and by connecting two analog CAM cells on the same MAL to represent one macro-cell, it is only possible to double the number of levels (i.e. $2^{N_{bits}}$) rather than the number of bits. Thus given memristors with $M=4$ bits, 16 analog CAM cells are needed for representing $N=8$ bits, leading to an exponential overhead. Here, we propose a novel solution to increase the precision with the same base analog CAM cell circuit, which requires 2 analog CAM cells and 2 clock cycles for performing 8-bit search operation with $M=4$ bits memristors.
To demonstrate, let us first take the case where the lower bound threshold $T_{L}$ is needed to be mapped and operated with $N=8$ bit of precision but we only have $M=4$ bits of precision in our analog CAM and memristor devices. The lower bound threshold can be sub-divided into its most significant bit representation $T_{LMSB}$ and least significant bit representation $T_{LLSB}$, each one made of $M$ bits and written as $T_{L}=2^{M}T_{LMSB}+T_{LLSB}=16T_{LMSB}+T_{LLSB}$. Each analog CAM cell computes the conditional operation $\textrm{MAL}=q\geq T_{L}$, where $q$ is the input query applied to the DL. We similarly divide the input $q$ into most and least significant representations $q_{MSB}$ and $q_{LSB}$, such that $q=2^{M}q_{MSB}+q_{LSB}=16q_{MSB}+q_{LSB}$ and we can obtain the result of MAL by computing
$\begin{split}[(q_{MSB}\geq T_{LMSB})\wedge(q_{LSB}\geq T_{LLSB})]\\ \vee(q_{MSB}\geq T_{LMSB}+1)\end{split}$ (2)
or alternatively, rearranging the terms:
$\begin{split}[(q_{MSB}\geq T_{LMSB}+1)\vee(q_{LSB}\geq T_{LLSB})]\\ \wedge(q_{MSB}\geq T_{LMSB}).\end{split}$ (3)
Note that Eq. 3 is more CAM-friendly since it is a conjunction of two terms that can be appended to the same MAL. The same calculation derives the upper bound $T_{H}$ conditions and with an AND operation between this and equation (3), the full computation $\textrm{MAL}=T_{L}\leq q<T_{H}$ results in:
$\begin{split}[(q_{MSB}\geq T_{LMSB}+1)\vee(q_{LSB}\geq T_{LLSB})]\\ \wedge(q_{MSB}\geq T_{LMSB})\\ \wedge[(q_{MSB}<T_{HMSB})\vee(q_{LSB}<T_{HLSB})]\\ \wedge(q_{MSB}<T_{HMSB}+1).\end{split}$ (4)
[…]
TABLE I: Input to apply for doubling precision with time shifting
Input Cycle 1 Cycle 2
$q_{HLSB}$ $q_{LSB}$ $GND$
$q_{LLSB}$ $q_{LSB}$ $VDD$
$q_{HMSB}$ $q_{MSB}$ $q_{MSB}-1$
$q_{LMSB}$ $q_{MSB}-1$ $q_{MSB}$
Table I shows the required inputs for each DL analog bus wire to compute equation (4) in a 2-step search operation. With MAL pre-charged high, the first cycle evaluates the first bracket of equation (4) for the lower and upper bound (the OR operation). Without pre-charging MAL again, the second cycle evaluates the second term of each bound by inputting always care (i.e. always mismatch) values to the LSB sub-cell. Therefore, an AND operation between the outputs of cycle 1 and cycle 2 is computed on MAL following cycle 2, similar to a 2-step search approach previously demonstrated for increasing bit density in TCAMs [30].
To the best of our knowledge, this is an entirely new approach and the proposed circuit operation is the most compact way of computing equation (4) with an analog CAM, and thus for doubling the bit precision. While other circuit designs may implement the same function in one clock cycle, these would have a significantly larger area. For example, one could separate cells performing the OR operation by connecting the discharge transistors in series rather than in parallel. This approach would also allow for selecting the precision of each feature to be either 4 or 8-bit, depending on the required precision. However, to perform such an operation in one cycle, three analog CAM cells with complex routing are needed. Given that analog CAM search latency is forecast to 100 ps [14], we found that performing two searches rather than one would not unduly increase the overall latency. Thus, we adopt this circuit formulation and operation which doubles the bit precision while doubling the area and analog CAM search latency.
III-C Core pipeline
In our ML experiments and hyperparameters search (see IV-A for details) we found out that $N_{leaves}=256$ per tree is typically sufficient to achieve high accuracy. Given that previous results have shown that analog CAM arrays with up to 128 rows do not encounter significant technology issues at 16 nm[14], each X-TIME core contains 2 stacked arrays, resulting in $2\times 128=256$ addressable words ($N_{leaves}$). We perform a survey on the size of typical tabular datasets [1], finding that usually $N_{feat}<100$ and often feature pre-processing reduces the number of $N_{feat}$ significantly before applying inputs to the decision tree ensemble. Previous results have shown that arrays with up to 100 columns have a very small access latency ($<$ 1ns) [14]. Nevertheless, we include an outlier dataset with $N_{feat}=130$[31] to show how to handle larger input vector sizes with input vector segmentation, as will be described below. Each core contains 2 arrays of 65 columns, resulting in a width of $2\times 65=130$. Overall, each core has a total addressable analog CAM of macro array size $2\times 128\times 2\times 65=256\times 65$.
The search operation is performed in series by groups of 65 features, while all the branches mapped to each queued array are computed in parallel. Multiple trees can be mapped to a single core, by arranging them in different portions of the analog CAM macro on the vertical axis. To avoid structural hazards, different considerations are made before applying new inputs to the core, as will be discussed below. When a feature vector is searched, a set of components with indices ranging from 00 to $64$ is sent to the first analog CAM logical array (aCAM1) while a second set of components with indices from $65$ to $129$ is sent to the second array (aCAM2). Comparison sequences on each tree branch are reconstructed by the AND operations made by aCAM1 and aCAM2 by selectively pre-charging only aCAM2 rows matched in aCAM1. Each analog CAM array has a latency of $\lambda_{CAM}=4$ cycles for performing the search operation. One cycle is used for pre-charging the MAL, two cycles for the search operation (MSB and LSB), and one cycle for latching the MAL sense amp, which leads to a total core latency of $\lambda_{C}=12$ cycles, with all other components (Buffer, MMR, SRAM, ACC) having a single cycle latency.
In the case of up to four trees mapped to the core when an array finishes the search operation, it starts processing a new sample and no additional bubbles are required. The maximum ideal core throughput $\tau_{C}$, considering a clock frequency of 1GHz [14] and a relatively large number of samples $N_{s}$, can thus be computed by
$\tau_{C}=\dfrac{N_{s}}{\lambda_{C}+\lambda_{CAM}(N_{s}-1)}\sim 250MSamples/s.$ (5)
If more than 4 trees are mapped in the core, extra pipeline bubbles $N_{B}=N_{trees,core}$ should be inserted. In fact, the MMR needs to perform more than 4 iterations in order to generate the one-hot vector for accessing the leaves, stored in the SRAM. The core pipeline in this case reaches an ideal throughput of
$\tau_{C}=\dfrac{N_{s}}{\lambda_{C}+N_{B}(N_{s}-1)}\sim 200MSamples/s\$ (6)
To maintain such a high throughput, cores need to be fed with input data fast enough. On the other hand, the architecture should be able to adapt to multiple models (such as RF, XGBoost, and others, either classification or regression problems) and exploit the vast amount of cores for parallel computations. Thus we designed a programmable NoC able to handle multiple workloads.
III-D Flexible connectivity
We identified four possible examples of inference operation that cover all ensemble models and tasks previously described, and exploit all the possibilities of our architecture: (1) regression and binary classification, (2) multi-class classification, (3) regression and binary classification with input batching and (4) multi-class classification with input batching.
In the case of regression and binary classification, leaves contain normalized partial predictions of each tree branch called logits, that are summed over the ensemble. In each core, the SRAM is programmed with logits that are summed locally in the core and in each network node inside the router. The final prediction is then sent to the CP which compares it with a threshold, typically 0 or 0.5, and sends the final result to the output. In the case of multiclass classification, logits should be summed class-wise and $N_{class}$ logits should be sent to the CP which performs an argmax computation to select the predicted class.
Figure 6: Example of network configurations for regression/binary classification (a), multiclass classification (b) and regression/binary classification with input batching (c). (d) Schematic representation of X-TIME programming flow
We propose a flexible/re-configurable NoC design (Fig. 6(a-c)) to aggregate results based on model type. The inset of Fig. 6(a) shows the downstream circuit schematic for the custom router. A buffer stores incoming leaves and has the output connected to a de-multiplexer (DEMUX) controlled by a configuration register. The DEMUX sends each leaf either to the accumulator circuit or directly to the output, with only logits belonging to the same class accumulated. Three examples of network configuration are shown in Fig. 6 for supporting (a) regression or binary classification, (b) multi-class classification, and (c) regression or binary classification with input batching. In the first case, the X-TIME compiler maps $N_{trees}=8$ trees with a depth $D=7$ each, with $N_{trees,core}=2$ each core. The same inputs (blue solid and light blue dashed lines) are applied to all the cores in parallel and all the logits (solid purple lines) are summed together, being a regression/binary classification problem. The X-TIME compiler sets the configuration bit of all routers to ’1’ such that logits are always accumulated in the router before propagating.
For multi-class classification (Fig. 6(b)), the X-TIME compiler maps $N_{estimators}=2$ estimators each made of $N_{trees}=4$ (one for each class) with a depth $D=7$. Each core holds two trees belonging to the same class and the same inputs (blue solid and light blue dashed lines) are applied to all cores in parallel. The X-TIME compiler sets router configuration bits to 0, as it is not possible to perform an accumulation/reduction at the router level and the reduction must be done globally by the CP. Thus, logits for each class are transmitted independently (solid purple lines, dashed purple lines, solid pink lines, and dashed pink lines).
Finally, for regression/binary classification with input batching (Fig. 6(c)), the X-TIME compiler maps an ensemble of $N_{trees}=4$ with depth $d=7$ and to increase the throughput, replicates trees in multiple cores, such that cores C1 and C2 map the same trees of cores C3 and C4. Different inputs are applied to different cores, or groups of cores by appropriately addressing each feature vector. Core C1 and C2 receive a first input (solid blue line) while cores C3 and C4 receive a second input (dashed light blue). Logits can be accumulated at the router level only between trees belonging to the same input batch, as in the lower level of the architecture in Fig. 6(c). The X-TIME compiler programs the lower level router configuration bits to ’1’ to enable local accumulation, and the higher level configuration bits to ’0’ to enable logits direct propagation. Input batching addresses one of the main issues of specialized accelerators in keeping resource utilization high [32]. With a similar procedure, multiple unique models can be mapped to the accelerator, by assigning a different batch to each model.
IV Evaluation Methodology
IV-A Datasets and model training
TABLE II: Datasets and models characterization
Dataset ID Task Samples $\boldsymbol{N_{feat}}$ $\boldsymbol{N_{classes}}$ Model $\boldsymbol{N_{trees}}$ $\boldsymbol{N_{leaves,max}}$
Churn modelling[33] 1 Binary Classification 10000 10 2 CatBoost 404 256
Eye movements[34] 2 Multiclass Classification 10936 26 3 XGBoost 2352 256
Forest cover type[35] 3 Multiclass Classification 581012 54 7 XGBoost 1351 231
Gas concentration[31] 4 Multiclass Classification 13910 129 6 Random Forest 1356 217
Gesture phase segmentation[36] 5 Multiclass Classification 9873 32 5 XGBoost 1895 256
Telco customer churn[37] 6 Binary Classification 7032 19 2 XGBoost 159 4
Rossmann stores sales[38] 7 Regression 610253 29 N/A XGBoost 2017 256
For our experiments, we trained state-of-the-art models with three ML algorithms on a set of benchmark datasets (details in Table II). The algorithms included random forests (RF) from scikit-learn [28], and gradient boosted trees from XGBoost [17], and CatBoost [18]. Fig. 6(d) shows a conceptual schematic of the workflow, whose output is a series of threshold maps (one for each tree) with the lower/upper bound to program in analog CAMs, the corresponding core/class address and leaf value for each node.
IV-B Cycle-detailed simulator
We developed a functional and cycle-detailed simulator [^1] https://github.com/HewlettPackard/X-TIME to evaluate latency, throughput, energy efficiency, area, and accuracy of the in-memory engine for tree-based model inference. The architecture simulator was developed using the Structural Simulation Toolkit (SST) [39] and estimates the performance on the benchmark datasets detailed above. SST, which is a parallel discrete-event simulator framework, can integrate modular components in a unified interface and simulate large-scale systems with MPI. It also provides interfaces for simulation components and components that perform actual functions such as processors, memory, and networks. For our system, the user-defined components emulate the cycle-level function of each block described in Fig. 4, and data flow through the links connecting the components with specified delay. The cycle-level architecture simulation consisting of an H-tree NoC with 4096 cores and 1365 routers verified the error-free functional operation and the pipelined architecture without hazards, while also yielding system performance estimates.
Figure 7: Area (a) and peak power (b) breakdown
Fig. 7 summarizes the power and area for the proposed architecture. Previous work generated layouts for the analog CAM at 16 nm  [14, 15], which included the analog CAM array, DAC, SA, and P-Ch area, and extracted power consumption using single cell parameters [15]. The analog CAM model was previously validated by experimental measurements [14]. Here, we simulate model mapping using memristor conductances from $1\mu S$ to $100\mu S$. The analog CAM model also includes parasitics that limit the CAM pre-charge and discharge time and the DAC operational frequency [14]. Power and area values for register and logic are based on TSMC 16nm PDK. The MMR design is based on the matching token MMR topology [40] whose output is fed back to MAL registers enabling multiple iterations when multiple trees are mapped to the same core. Notably, the analog CAM arrays dominate area and power consumption, with peripheral components contributing negligibly. Also, the overall peak power consumption of 19 W is comparable to GPU idle power ($\sim$25 W).
IV-C Profiling Tree Model Inference on NVIDIA GPUs, FPGA and SRAM based accelerators
Figure 8: Schematic of the SRAM-based accelerator core used as digital baseline.
To demonstrate the analog CAM performance capabilities, we compare our results to three alternative accelerator approaches: GPU, FPGA, and digital SRAM. As our first performance baseline, we profiled the inference step of the tree models and corresponding datasets shown in Table II on an NVIDIA V100 GPU. We measured each GPU kernel execution time with 5 repetitions to process data batches of increasing size, up to a saturation point, where throughput improvements were no longer observed. To estimate GPU upper bound performance, we isolated the kernel execution time of each model using NVIDIA’s nvprof tool, and did not include the memory transfer times between host (CPU) to device (GPU) or internal GPU transfers in our measured times. When possible, we profiled models using CUDA kernels from the NVIDIA RAPIDS Forest Inference Library (FIL).
Moreover, we compared X-TIME with an FPGA baseline FAXID [8]. FAXID is a custom hardware accelerator for low-latency ransomware detection using high-level synthesis (HLS) using binary classification models. As FAXID only supports binary classification, for our comparison we used a 1-vs-all approach to benchmark latency and throughput where we select one of the classes as class 0 and set the remaining classes to class 1. In FAXID the maximum depth was restricted to 6, as their ransomware detection accuracy did not increase with deeper trees. To keep the baseline results faithful to the FAXID HLS design, we optimized the hyperparameters of Table II for smaller depth and larger number of trees on FAXID to match accuracy with the other approaches.
Finally, we compare X-TIME to a custom digital accelerator baseline using SRAM to isolate the performance benefits from the analog CAM specifically. This SRAM baseline uses a distributed architecture with each tree mapped to a small SRAM accessed serially by a floating point unit [5], as shown in Fig. 8. Each SRAM word stores the node value and feature, the left and right children addresses, or the corresponding logit or class in the case of a leaf node. We considered the X-TIME network architecture and changed just the core operation (analog CAM or digital SRAM) to simulate the effect on throughput, of multiple accesses required for each tree.
One might consider a digital TCAM to be another meaningful comparison. However, as previously demonstrated [41], using a traditional digital TCAM for performing on-access comparison would result in large overheads. In detail, $2^{8}$ TCAM columns would be needed to represent each feature with 8-bit precision, and thus tree inference of 129 features requires 33024 TCAM columns. Considering an 80-column TCAM macro has an access time of $\sim 500ps$ [42], inference with a TCAM would take $500\cdot 33024/80ps\approx 20us$ per tree, significantly larger than the few cycles analog CAM latency.
V Results
V-A Model inference performance
Figure 9: (a) Accuracy for different training constraints and (b) relative accuracy as a function of memristor and DAC defects.
In designing our HW architecture, we investigated the required bit accuracy and model types that the architecture should support to maintain state-of-the-art model accuracy. To assess these parameters, we compared model accuracy for different training parameters and model types to guide the X-TIME requirements (Fig. 9(a)). First, we set our baseline, achieving state-of-the-art accuracy [43] by training all the models without constraints in the hyperparameters search (’Unconstrained’ in Fig. 9(a)). We then optimized models to fit the hardware using the proposed 8-bit precision ACAM (‘X-TIME 8bit’: maximum $N_{trees}=4096$, $N_{leaves,max}=256,N_{bit}=8$), which resulted in limited accuracy loss compared to the unconstrained baseline case. To assess the need for 8-bit vs. 4-bit analog CAMs, we optimized the model to fit hardware using a 4-bit ACAM [15], allowing the number of leaves to double, such that the area would be equivalent to ’X-time 8bit’(‘X-TIME 4bit’: maximum $N_{trees}=4096$, $N_{leaves,max}=512,N_{bit}=4$). Another option would double the maximum number of allowed trees, but we generally found that 4096 trees are high enough for matching unconstrained baseline accuracy. Our results showed a reduction in performance of $20\%$ in the case of regression on Rossmann store sales dataset [38] and $18\%$ in the case of multi-class classification on gas concentration dataset [31]. Finally, we limited the optimizer to RF models only, which resulted in significant accuracy degradation for several datasets (label = ‘Only RF’). Given that it is not possible to train directly with 4-bit precision, the after-training quantization significantly decreased accuracy, demonstrating the need for supporting a broader range of model types beyond those demonstrated in previous works [44, 15].
To assess the promise of this analog hardware approach in practical use, we also studied the impact of variations and defects of the analog hardware on model inference accuracy. The two most impactful sources of error to classification/regression results in this analog hardware are the known memristor conductance variations and DAC input variations to the aCAM. We considered experimental data from TaOx memristor and 16 nm analog circuits[15]. Fig. 9(b) shows the mean relative accuracy, defined as the ideal accuracy of Fig. 9(a) over the defect-compromised accuracy as a function of the percentage of defects. We defined a defect as a 1-level random flip in the memristor conductance or DAC output voltage.
The plot shows the average results across all classification datasets. We observe that these models and our architecture are fairly resilient to memristor variations, as individual errors do not have a large impact due to the ensemble nature of the model itself. With the aforementioned conductance window and a conservative standard deviation of $1\mu S$ it is expected to have a flip probability of $\sim 0.2\%$, corresponding to accuracy drop $<0.5\%$. On the other hand, while DAC variations also impact the accuracy, DACs are usually designed for operating up to 16 bits [45], thus realizing a stable/error-free 4-bit design is not particularly challenging.
V-B Comparison with state-of-the-art
Figure 10: Comparison of (a) latency and (b) throughput of this work, GPU, FPGA, and SRAM based accelerator from Fig. 8 (a) Throughput as a function of $N_{trees}$ and $D$.
We benchmarked our proposed architecture on a wide range of models (corresponding to X-TIME 8bit of Fig. 9) and dataset types (Table II), and compared latency and throughput (see Fig. 10(a,b)) to GPU, FPGA and a digital accelerator based on small SRAMs and similar to Booster [5]. When possible, we applied input batching and replicated trees to increase throughput. As shown, our architecture produces drastically reduced latency, frequently to $\sim$100ns compared to GPU at 10 $\mu$s to ms, while also demonstrating significantly improved throughput by 10-120$\times$ across most datasets. Our architecture shines for large binary classification or regression models with a peak improvement in latency by 9740×9740$\times$9740 × at 119$\times$ higher throughput for the Churn modeling dataset inference[33] compared to GPU. As tree search is parallel in each analog CAM array of each core, as well as the in-network reduction, latency, and throughput are constant with $N_{trees}$ and $N_{leaves}$ with a delay solely dependant on $N_{feat}$ and $N_{classes}$. In contrast, on GPU there is a linear dependence on $N_{trees}$ and $D$, which limits the performance for large models. In the FPGA implementation, namely FAXID, there is a linear dependence on $N_{trees}$. The SRAM-based accelerator has a linear dependence on $D$ (since multiple memory accesses are required) and $N_{trees}$ when more than a tree needs to be mapped in a core. As an example, we studied the throughput as a function of $N_{trees}$ and $D$ for GPU, FPGA, SRAM-based accelerator, and X-TIME as shown in Fig. 10(c), for a multiclass classification dataset [35], demonstrating the predicted trend. We note that the SRAM-based accelerator still suffers from load imbalance between trees, similar to GPU, in which trees with different depths have different inference times, which results in the need for synchronization before the reduction operation. The high throughput of X-TIME can produce an extremely low energy per decision, down to 0.3 nJ/Dec, highlighting the potential for energy-efficient applications.
VI Conclusions
In this work, we presented a novel architecture for accelerated tree-based ML inference with in-memory computing based on a novel increased precision analog non-volatile CAMs. Our approach maps decision trees to analog CAM ranges, which compute directly the matched results in parallel given an input feature vector. Leaf values are stored in a local SRAM and accumulated at the core level when multiple trees are mapped in the same core. Outside the core, an H-tree NoC connects 4096 cores and a co-processor in a single chip. An in-network computing structure accumulates leaf values as data flows to limit data movement. The on-chip network is re-configurable to map different models (regression, multi-class classification, etc) and inference techniques, such as input batching. Our results show significant improvement in latency – by almost four orders of magnitude, compared with an NVIDIA V100 GPU, at a throughput up to 120$\sim$ 120$\times$ 120 × higher across a wide range of dataset types and model topologies.

### v4

@@
-X-TIME: An in-memory engine for accelerating machine learning on tabular data with CAMs
+
 Abstract
@@
 Previous work has described the use of an analog content addressable memory (CAM) component for efficiently mapping random forests.
-In this work, we focus on an overall analog-digital architecture implementing a novel increased precision analog CAM and a programmable network on chip allowing the inference of state-of-the-art tree-based ML models, such as XGBoost, CatBoost, and others.
-Results evaluated in a single chip at 16nm technology show 119$\times$ higher throughput at 9740$\times$ lower latency compared with a state-of-the-art GPU, with a 19W peak power consumption.
-Index Terms:
-In-memory computing, tree-based ML, content addressable memories
-I Introduction
-Extracting relevant information from structured (i.e., tabular) data is paramount in practical data science.
-Tabular data consist of a matrix of samples organized in rows, each of which has the same set of features in its columns, and is commonly used for medical, financial, scientific, and networking applications, enabling efficient search operations on large-scale datasets.
-While deep learning has shown impressive improvements in accuracy for unstructured data, such as image recognition or machine translation, tree-based machine learning (ML) models still outperform neural networks in classification and regression tasks on tabular data [1].
-This difference in model accuracy is due to robustness to uninformative features of tree-based models, bias towards a smooth solution of neural networks models, and in general the need for rotation invariant models where the data is rotation invariant itself, as in the case of tabular data [1].
+In this work, we develop an analog-digital architecture that implements a novel increased precision analog CAM and a programmable chip for inference of state-of-the-art tree-based ML models, such as XGBoost, CatBoost, and others.
+Thanks to hardware-aware training, X-TIME reaches state-of-the-art accuracy and 119$\times$ higher throughput at 9740$\times$ lower latency with >150×>150$\times$> 150 × improved energy efficiency compared with a state-of-the-art GPU for models with up to 4096 trees and depth of 8, with a 19W peak power consumption.
+{IEEEkeywords}
+Content Addressable Memories, Decision Trees, In-memory Computing, ReRAM
+1 INTRODUCTION
+[…]
+Extracting relevant information from structured (i.e., tabular) data is the foundation of practical data science.
+Tabular data consists of a matrix of samples organized in rows, each of which has the same set of features in its columns, and is commonly used for medical, financial, scientific, and networking applications, enabling efficient search operations on large-scale datasets, making them one of the most common sources of data in the data science industry.
+Moreover, data collected in real-time for heterogenous sources, for example during scientific experiments, are organized in tabular data and with fast processing may lead to in-the-loop intelligent control.
+While deep learning has shown impressive improvements in accuracy for unstructured data, such as image recognition or machine translation, tree-based machine learning (ML) models still outperform neural networks in classification and regression tasks on tabular data [1], including modern Transformers for tabular data [2].
+This difference in model accuracy is due to robustness to uninformative features of tree-based models, bias towards a smooth solution of neural networks models, and in general the need for rotation variant models where the data is rotation variant itself, as in the case of tabular data [1].
 With high model accuracy, ease of use, and explainability [3] tree-based ML models are greatly preferred by the data science community [4], with $>74\%$ of data scientists preferring to use these models while $<40\%$ choose neural networks in a recent survey.
@@
 While small models are easily executed, performance issues for larger tree-based ML models arise due to limited inference latency both on CPU and GPUs [5, 6] originating from irregular memory accesses and thread synchronization issues.
-As large tree-based models are finding performance limits in existing hardware, larger models ($>$1M nodes) are demonstrating increasingly compelling uses in practical environments where high throughput and moderate latency are desired.
+As large tree-based models ($>$1M nodes) are finding performance limits in existing hardware, they are demonstrating increasingly compelling uses in practical environments where high throughput and moderate latency are desired.
 For example, the winner of a recent IEEE-CIS Fraud detection competition used a tree-based ML model with 20M nodes [7].
@@
 This approach eliminates access, transfer, and processing of data in a separate compute unit [11], reducing latency.
-Previous research has shown that these devices can efficiently perform multiplication functions in crossbar architectures [11].
+Previous research has shown that multiplication functions can efficiently be performed in crossbar architectures [12].
 However, crossbar architectures are limited to matrix operations, which are not useful for inference on tree-based models.
 Another class of in-memory computing primitives is Content Addressable Memories (CAMs) [13, 14], high-speed and highly parallel circuits that search for a given input in a table of stored words and return the matched search input’s location.
-More recently, analog CAMs have been proposed [14] which leverage the analog states in memristor devices.
-An analog CAM cell performs in situ comparisons of stored (analog) ranges with (analog) inputs, by computing if the input is within the range.
-Effectively, analog CAM arrays perform a massively parallel comparison between an input vector of analog values and stored words of analog ranges.
-Given that DTs are a set of conditional branches, analog CAMs naturally map DTs and perform highly parallel inference of decision forests [15].
-Previous work showed the potential of this approach but was limited to component-level estimation, 4-bit precision, and random forest models [15].
+CAMs use relatively cheap peripherals, e.g.
+sense amplifier circuits rather than expensive ADCs required by crossbar arrays for in-memory computing operations.
+More recently, analog CAMs have been proposed [15] which leverage the analog states in resistive switching memories (RRAM) [16].
+An analog CAM performs the comparisons of stored (analog) ranges with (analog) inputs, by computing if the input is within the range.
+Given that DTs are a set of conditional branches, analog CAMs naturally map DTs and perform highly parallel inference of decision forests [17, 18].
+Our previous work showed the potential of this approach but was limited to component-level estimation, 4-bit precision, and random forest models [17], while the full architecture study in this work - which includes designs for higher precision Analog CAM operation, study of peripheral circuits and builds in support for multiple ML models convincingly demonstrates the feasibility of the initial concept.
 This work presents X-TIME, an in-memory engine for accelerating tree-based ML models with analog CAMs.
-In this work we demonstrate the need for a holistic architecture design, 8-bit operation, and support for multiple tree-based ML models, overcoming the limitation of previous work [15].
-The major contributions of this paper compared to previous work [15] are (1) a novel analog CAM cell design and operation to perform decision tree inference at doubled (i.e.
-8-bit) memristor precision (i.e.
-4-bit), (2) an in-core and in-network leave accumulation architecture to seamlessly perform inference of a large number of trees in parallel, (3) a programmable Network-on-Chip (NoC) to operate any tree-based ML model regardless of the ensemble reduction operation, (4) a compiler supporting the mapping of multiple tree-based ML models and (5) a rigorous characterization method and scaling comparison with respect to GPUs, FPGAs, and conventional digital accelerators.
-Our simulated evaluation shows up to 119$\times$ increased throughput at 9740$\times$ decreased latency in a single 19W chip.
-II Background
+In this work we demonstrate the need for a vertical architecture design, 8-bit operation, and support for multiple tree-based ML models, overcoming the limitations of previous work [17].
+The major contributions of this paper compared to previous work [17] are at the circuit level(1) a novel analog CAM cell design and operation to perform decision tree inference at doubled (i.e.
+8-bit) resistive memory precision (i.e.
+4-bit), (2) a complete design of the core and network operation, and at the architecture level (3) an end-to-end design targeting a large scale accelerator, (4) a compiler (X-TIME-C) supporting the mapping of multiple tree-based ML models to the HW starting from established intermediate representation [19], and at the co-design level (5) an HW-aware model training algorithm for reaching state of the art accuracy and (6) a complete simulation framework to benchmark and rigorous characterization comparison against NVIDIA V100 GPUs, FPGAs, and conventional digital accelerators.
+Our simulated evaluation shows up to 119$\times$ increased throughput at 9740$\times$ decreased latency with >150×>150$\times$> 150 × improved energy efficiency in a single 19W chip.
+2 BACKGROUND
 Figure 1: Schematic illustration of (a) a decision tree and (b) a decision tree ensemble.
-II-A Tree-based Machine Learning
+2.1 Tree-based Machine Learning
 A DT is a regression and classification ML model consisting of a set of nodes arranged in a tree-like structure.
@@
 In the case of RF, the reduction operation consists of a majority voting, while in all the others leaf values of each DT are summed.
-II-B Tree-based ML model inference on GPU
-While training DT models is relatively fast when compared to, for example, a large deep neural network, the computations involved in inference are very irregular and not particularly suitable for CPUs and GPUs, becoming inefficient for very large models, with $>1$ M nodes [6].
+While training DT models is relatively fast when compared to, for example, a large deep neural network, the computations involved in inference are very irregular and not particularly suitable for CPUs and GPUs [6], becoming inefficient for very large models, with $>1$ M nodes [7].
 While CPUs can run irregular control paths fairly well, their parallelism is limited.
-GPUs have a very large parallelism, but suffer from the irregularity of memory access patterns, as will be explained in detail below.
-Recent NVIDIA libraries implementations have shown order of magnitude improvements compared to CPU and previous GPU implementations  [19].
-However, three different factors still limit GPU throughput and latency [4], in particular for tree ensembles.
-First, a high number of uncoalesced memory accesses, due to the low ratio between requested and fetched data, causes low memory efficiency.
-Trees are usually stored in memory with nodes from different trees interleaved in breadth-first order, resulting in coalesced memory accesses close to the root, where children nodes are more likely to be part of the same memory transaction of the parent node, but uncoalesced memory accesses further from the root.
-Thus, with increasing tree depth the proportion of coalesced memory accesses decreases and memory system efficiency also decreases.
-Second, synchronization between threads and the consequent load imbalance due to waiting for the deepest trees limits latency.
-A common implementation scheme on GPU assigns each tree in the ensemble to a different thread by round-robin.
-At the end of each thread block, a reduction operation is then performed.
-However, a given ensemble may contain short and tall trees, with the latter having substantially larger latency.
-Thus, each thread block must wait until the slowest (tallest) tree has finished processing, increasing overall latency.
-Finally, with a larger number of trees $N_{trees}$ in the ensemble, the increasing number of thread blocks and the global reduction performed between thread blocks becomes a significant overhead.
-Multiple techniques have been proposed to increase the throughput of tree ensemble inference [19, 4], mainly by re-organizing the memory - for example, for sparse trees and scheduling of threads - but with limited impact on latency.
-In this work, we focus on addressing these limitations with an in-memory computing architecture using content-addressable memories.
+GPUs have a very large parallelism, but suffer from the irregularity of memory access patterns, as explained in detail in Supplementary Information LABEL:supp-si:section:gpu_inference.
 Figure 2: Comparison of (a) a conventional CAM/TCAM and (b) an analog CAM
-II-C Analog Content Addressable Memory
+2.2 Analog Content Addressable Memory
 CAMs are specialized memory devices that search a table of stored words (rows) or keys in parallel for a given input query.
-The CAM search output is a one-hot vector where a high value corresponds to a match between the input data query and stored row word key [12, 13].
+CAMs output a one-hot vector where a high value corresponds to a match between the input data query and stored row word key [13, 14].
 In conventional CAMs, this output vector is converted into a single address by a priority encoder.
 Each CAM cell functions as both memory and bit comparison, and the input query is compared to each CAM word simultaneously to produce a high throughput, low latency few-cycle compare operation at the cost of high power and area.
-In addition to binary (0 and 1) values, ternary CAMs (TCAMs) have a third wildcard value, X (or “don’t care”), which can be used to compress the stored CAM table or perform partial match searches.
+In addition to binary (0 and 1) values, ternary CAMs (TCAMs) have a third wildcard value, X (or don’t care), which can be used to compress the stored CAM table or perform partial match searches.
 When an X is stored or searched, that cell matches irrespective of whether the searched or stored value is 1 or 0.
 Fig.
-2(a) shows a schematic illustration of a conventional CAM/TCAM circuit, where input queries are applied on the column search lines (SLs) and outputs are read along the row ”match lines” (MALs).
-An XNOR operation occurs between the query values and stored cell values in the corresponding column.
+2(a) shows a schematic illustration of a conventional CAM/TCAM circuit, where input queries are applied on the column Search Lines (SLs) and outputs are read along the row Match Lines (MALs).
+Essentially, an XNOR operation occurs between the query values and stored cell values in the corresponding column.
 The MAL $j_{th}$ row output is produced by a bit-wise AND operation between cells along the row and returns a match (high) or mismatch (low).
@@
 While offering high performance, conventional SRAM-TCAMs are power and area-hungry and require 16 transistors.
-Recently, there has been significant interest in developing TCAMs with memristor devices and resistive non-volatile memories [20, 21, 22] with compact size and low-power operation.
+Recently, there has been significant interest in developing TCAMs with RRAM devices  [23, 24, 25] with compact size and low-power operation.
 As programming these devices is relatively slow, emerging TCAMs have targeted in-memory computing applications [26, 14] where stored value updates are infrequent.
-Beyond direct digital memory operation, memristor devices have also demonstrated ranges of stable analog states of $>16$ levels.
+Beyond direct digital memory operation, RRAM devices have also demonstrated ranges of stable analog states of $>16$ levels [27].
 Leveraging this analog state capability, recent work proposed and demonstrated an analog CAM circuit design and operation [15] (conceptual representation in Fig 2(b)).
@@
 $\text{MAL}_{i}=\bigwedge_{j}{(T_{L,ij}\leq q_{j}<T_{H,ij})}\text{,}$ (1)
-where $q$ is the query input vector applied to DL, and $T_{L}$ and $T_{H}$ are the stored range thresholds for the CAM cell range lower and upper bounds, respectively.
+where $q$ is the query input vector applied to Data Line (DL), and $T_{L}$ and $T_{H}$ are the stored range thresholds for the CAM cell range lower and upper bounds, respectively.
 In the analog CAM circuit, no analog-to-digital conversion is required, eliminating an expensive component which is a limiting bottleneck of other analog in-memory computing primitives [28].
-Multiple analog CAM circuits have been realized using different technologies such as memristor devices [14] and ferroelectric memory devices [25].
+Multiple analog CAM circuits have been realized using different technologies such as RRAM devices [15] and ferroelectric memory devices [29].
 In such circuits, two analog memory elements represent the lower and upper bound thresholds.
@@
 Figure 3: DT example from Fig.1(a) with nodes mapped in an analog CAM and leaves mapped into a RAM.
-II-D Mapping tree models to analog CAMs
-The X-TIME compiler takes as input the description of a DT ensemble in a tabular mode, where each row contains the node attributes (feature ID and value) and the left/right children.
-The compiler traverses the tree structures extracting all the root-to-leaf paths.
-Finally, it maps each of them to an ACAM row.
-As an example, Fig.
-3 shows the mapping of Fig.
-1(a).
-Each column corresponds to a different feature of the input vector $q$.
+2.3 Mapping tree models to analog CAMs
+Fig.
+3 shows the mapping of the decision tree in Fig.
+1(a) to the Analog CAM.
+Each row represents a branch of the tree and each column corresponds to a different feature of the input vector $q$.
 Each node is mapped in the column corresponding to the feature used in the node’s comparison.
 The comparison’s threshold value is programmed in the analog CAM’s lower and/or upper bound, according to the direction, left or right, taken by the comparison’s result for the root-to-leaf path.
-A “don’t care” is used if a feature is missing, by programming the full range (0 to 1 in our example) in the analog CAM cell.
+A don’t care symbol is used if a feature is missing, by programming the full range (0 to 1 in our example) in the analog CAM cell.
 The feature vector (query) is applied on the DL and the output on the MAL is high for the matched branch, i.e.
@@
 This mapping requires the number of CAM rows to match the number of tree leaves, and the number of CAM columns to match the number of features $N_{feat}$.
-MALs can be connected directly to the Word Lines (WLs) of a conventional RAM storing leaf values to return the model’s prediction.
+MALs can be connected directly to the Word Lines (WLs) of a conventional SRAM storing leaf values to return the model’s prediction.
 Effectively, the analog CAM traverses root-to-leaf paths, and the RAM retrieves leaf values.
@@
 In addition, we extend the use of the previously presented analog CAM [15, 17] circuit to support higher precision data by developing a flexible connectivity strategy and bit-arithmetic refactoring.
-We compare the performance of our architecture on state-of-the-art tabular dataset ML tasks against GPU, FPGA and digital ASIC, showing a large improvement in latency and throughput while using a single chip with limited power consumption.
-III Architecture organization
+We compare the performance of our architecture on state-of-the-art tabular dataset ML tasks against GPU, FPGA, and digital ASIC, showing a large improvement in latency and throughput while using a single chip with limited power consumption.
+3 Hardware-aware model training
+To ensure a targeted and relevant accelerator design, we explore the HW algorithm co-design space by assessing model types, sizes, and accuracy requirements in conjunction with different analog CAM HW and architecture parameters.
+First, we consider the limitations of the previously presented RF accelerator based on analog CAM [17].
+The accelerator was limited to 4-bit precision, essentially due to the programming accuracy of ReRAMs, and it was benchmarked on a single dataset.
+However, different tree-based ML models would perform differently for different tasks, such as binary or multiclass classification or regression, and different datasets.
+The 4-bit operation might not be enough to capture more complex datasets that would require a larger number of bins for each feature.
+Finally, the maximum depth of each tree and the number of trees, corresponding to the number of words in the analog CAM and the number of analog CAM arrays, should be optimized on a variety of datasets to ensure a general well-performing accelerator.
+Thus we performed an in-depth characterization of tree-based ML models, namely RF, XGBoost, and CatBoost, performance on several datasets to understand (1) the necessity for mapping diverse models to the accelerator, (2) the size of each analog CAM macro and total number of macros in the accelerator and (3) the required bit precision for the analog CAM operation.
+Table 1: Datasets and models characterization
+Dataset ID Task Samples $\boldsymbol{N_{feat}}$ $\boldsymbol{N_{classes}}$ Model $\boldsymbol{N_{trees}}$ $\boldsymbol{N_{leaves,max}}$
+Churn modelling[32] 1 Binary Classification 10000 10 2 CatBoost 404 256
+Eye movements[33] 2 Multiclass Classification 10936 26 3 XGBoost 2352 256
+Forest cover type[34] 3 Multiclass Classification 581012 54 7 XGBoost 1351 231
+Gas concentration[35] 4 Multiclass Classification 13910 129 6 Random Forest 1356 217
+Gesture phase segmentation[36] 5 Multiclass Classification 9873 32 5 XGBoost 1895 256
+Telco customer churn[37] 6 Binary Classification 7032 19 2 XGBoost 159 4
+Rossmann stores sales[38] 7 Regression 610253 29 N/A XGBoost 2017 256
+We perform a survey on the size of typical tabular datasets [1], finding that usually $N_{feat}<100$ and often feature pre-processing reduces the number of $N_{feat}$ significantly before applying inputs to the decision tree ensemble.
+Nevertheless, we include an outlier dataset with $N_{feat}=130$[35] to show how to handle larger input vector sizes with input vector segmentation, as will be described later.
+Table 1 shows a characterization of the selected datasets.
+Figure 4: (a) RMSE as a function of model size in the case of Rossmann stores sales[38] for multiple configurations.(b) Best accuracy of each dataset for multiple configurations.
+First, we trained all the models optimizing the hyperparameters without limitations, similar to what would happen if training the model for the GPU.
+Then, we studied the impact of bit precision.
+Given that from implementing 4-bit to 8-bit operation the analog CAM area would roughly double, as explained in Section 4.2, we doubled the number of available trees in that case.
+Finally, we limited the training to RF to demonstrate that it underperforms compared to other models, such as XGBoost, thus a general architecture that accommodates gradient-boosted models, as opposed to the previously realized accelerator [17] for RF-only is vital.
+Fig.
+4(a) shows as an example the Root Mean Square Error (RMSE) as a function of the model size expressed as the size of an analog CAM-based hardware, namely $N_{trees}\times 2^{D}$, with $D$ maximum depth and $N_{trees}$ number of trees in the case of Rossmann stores sales dataset[38].
+We repeated the same experiments for all datasets (Fig.
+LABEL:supp-si:fig:dse and LABEL:supp-si:fig:dse2), picking as design choices for our architecture $N_{trees}=4096$ and $D=8$, corresponding to a total of 4096 cores of with $2562^{8}=$ 256 words each.
+The dashed line corresponds to our selected maximum model size, and thus hardware size and light blue area correspond to the model that we can map to the hardware.
+It is possible to see how the accuracy saturates at a certain model size regardless of the configuration chosen (i.e.
+4 or 8-bit, RF-only, or any tree-based model).
+Results show that RF models are insufficient to reach good precision, and 8-bit inference can increase accuracy up to 2% for classification and a $\sim$20reduction on RMSE for regression.
+Fig.
+4(b) shows the accuracy and RMSE for multiple datasets assuming constraint-free hyperparameters optimization (GPU), limited to our design choice with 8-bit (X-TIME 8-bit) and 4-bit (X-TIME 4-bit), and using only RF models (RF Only) demonstrating that (1) 8-bit can match the constraint-free design, (2) 4-bit is not enough for some datasets and (3) RF is underperforming significantly the best models for some datasets.
+A printout of the data can be also found in Table  LABEL:supp-si:tab:accuracy.
+The final model and parameter selections for each dataset are reported in Table 1.
+Given that the hardware cost scales aggressively with $D$ but linearly with $N_{trees}$ (Fig.
+LABEL:supp-si:fig:cost), and the accuracy saturates for $D>8$, we designed X-TIME for being able to perform inference of models up to $N_{trees}=4096$, $D=8$, and $N_{bit}=8$.
+If more trees are needed, multiple X-TIME chips can be operated in parallel, similarly to how multiple cores are operated in parallel.
+4 X-TIME Architecture
 Figure 5: Architecture hierarchy
@@
 The accelerator is composed of a group of cores connected by an H-tree NoC topology converging to a co-processor (CP).
-Each core contains an analog CAM array acting as a processing element (PE-CAM), buffer, multiple match resolver (MMR), SRAM memory, and accumulator (ACC) (i.e.
-feedback adder).
-Typical CAM architectures also contain similar elements but the memory is used as an address encoder rather than for storing leaves [12].
-To map models larger than a single array, each PE-CAM has a number of stacked arrays (extended row-wise) and queued arrays (extended column-wise) to handle more and larger words, respectively.
+Each core contains an analog CAM array acting as a processing element (PE-CAM) and an SRAM memory.
+To map models larger than a single array, each PE-CAM has stacked arrays (extended row-wise) and queued arrays (extended column-wise) to handle more and larger words, respectively.
 Stacked arrays share the same peripherals, namely the Pre-Charger (P-Ch), the Digital Analog Converter (DAC), the Sense Amplifier (SA), and the Match-Line Register (ML-REG).
 The connection of queued arrays ensures that only previously matched lines are charged by making the ML-REG of array $i$ feed the P-Ch of the array $i+1$.
-III-A Inference
-X-TIME compiler maps the trained trees to the cores according to Fig.
+4.1 Inference
+The X-TIME Compiler (X-TIME-C) maps the trained trees to the cores according to Fig.
 3.
-Training is performed using standard libraries such as scikit-learn[28] or XGBoost[17].
-XGBoost outputs a table with $N_{leaves}+N_{nodes}$ (where $N_{leaves}$ is the total number of leaves in the ensemble) rows each storing the node ID, left and right child ID, node value, node feature, leaf value (if it is a leaf) and tree ID.
-Converting this table for CAM usage, we obtain a new table of size $L\times(2N_{feat}+3)$ with each row storing the lower/upper bound for each feature, the leaf value, class ID/label (in the case of multi-class classification) and tree ID, representing the model ensemble for our architecture.
-The X-TIME compiler assigns trees to cores via round-robin, directly mapping the first $L\times(2N_{feat})$ columns of a tree to the lower and upper bound of the analog CAM circuit, and the third-last column to the SRAM.
+X-TIME-C is based on Treelite [19], a universal model exchange and serialization format for tree-based models.
+Treelite outputs a C-like code with conditional statements representing each tree in the ensemble as an intermediate representation for further compilation to the hardware.
+X-TIME-C converts this code for analog CAM usage, obtaining a table of size $L\times(2N_{feat}+3)$ with each row storing the lower/upper bound for each feature, the leaf value, class ID/label (in the case of multi-class classification) and tree ID, representing the model ensemble for our architecture.
+X-TIME-C assigns trees to cores via round-robin, directly mapping the first $L\times(2N_{feat})$ columns of a tree to the lower and upper bound of the analog CAM circuit, and the third-last column to the SRAM.
 Class and tree ID are uniquely represented in the core address.
-The largest ensemble in our architecture is constrained by the maximum number of leaves $N_{leaves,max}$ in all trees and the number of available cores $N_{C}$.
-If $L<N_{words}/2$, multiple trees can be mapped to a single core, where $N_{words}=N_{stacked}\times H$ is the number of words addressable by each core (with $H$ number of rows in an analog CAM array and $N_{stacked}$ number of arrays stacked).
+Fig.
+LABEL:supp-si:fig:compiler shows a schematic visualization of X-TIME-C compilation flow.
+The largest ensemble that can be mapped to X-TIME is constrained by the maximum number of leaves $N_{leaves,max}$ in all trees and the number of available cores $N_{C}$.
 During inference, a feature vector is sent through the network to a core and searched in the PE-CAM.
 Each queued analog CAM block receives a different portion of the feature vector which is converted into analog voltage values by a DAC.
-A shift operation before the conversion is performed to increase operation precision (as discussed in section III-B).
-A logical AND operation is inherently performed between the queued arrays, similar to MAL segmentation used in modern TCAMs to maintain high throughput where MALs are divided into sections of shorter words with an AND between individual MAL results.
-Connecting multiple TCAM arrays with common MALs is logically equivalent to a larger TCAM, with the bit-wise AND operation between TCAM cells on the same MAL now extended between arrays.
-In our case, the analog CAM stores a branch split into multiple analog CAM arrays and reconstructed by means of the logical AND, such that only if all branch nodes arranged in multiple arrays are matched will a match be returned on the corresponding MAL.
-The final output of the PE-CAM is stored in a buffer and fed into the MMR which selects only one match signal to proceed logically.
-The MMR is needed due to the possibility of mapping multiple trees in a single core.
-In fact, given $N_{trees,core}$ mapped to a core, for each inference, the PE-CAM would return $N_{trees,core}$ matches in the form of a sparse Boolean vector with $N_{trees,core}$ match values (“1”s).
-The MMR generates $N_{trees,core}$ one-hot Boolean vectors used for directly accessing the SRAM WLs.
-The SRAM stores the corresponding leaf value of each branch, which is retrieved sequentially and accumulated by the ACC block, which performs a local reduction operation.
-After $N_{trees,core}$ accumulations are performed, the resulting leaf value is sent as output to the router.
-Leaf values coming from multiple cores can be summed at the router level as section III-D discusses, performing a shared reduction operation.
+A logical AND operation is inherently performed between the queued arrays, similar to ML segmentation used in modern TCAMs to maintain high throughput where MLs are divided into sections of shorter words with an AND between individual ML results.
+Connecting multiple TCAM arrays with common MLs is logically equivalent to a larger TCAM, with the bit-wise AND operation between TCAM cells on the same ML now extended between arrays.
+In our case, the analog CAM stores a branch split into multiple analog CAM arrays and reconstructed using the logical AND, such that only if all branch nodes arranged in multiple arrays are matched will a match be returned on the corresponding MAL.
+The final output of the PE-CAM is stored in a register and used to address the SRAM, storing the leaf value of each branch.
+Leaf values coming from multiple cores are summed with a programmable adder tree to perform the reduction operation.
 Finally, the resulting leaf value is propagated to the CP which performs a global reduction and computes the model prediction.
 The latter is usually a simple operation, such as a majority voting or a threshold comparison.
-III-B Analog CAM with improved precision
-Figure 5: (a) analog CAM 8-bit macro-cell and (b) analog CAM 4-bit sub-cell circuit.
-ML algorithms usually do not require high-precision floating point operations for achieving good results [29].
-As will be detailed in section IV we found out that typically 8 bits precision, corresponding to 256 bins for each feature, is required for matching floating point (FP) accuracy inference of tree-based ML models.
-While analog CAM circuits offer the possibility of mapping arbitrary ranges, memristor device reliability poses a challenge in supporting more than 16 levels or 4 bits [15].
-In the case of crossbar arrays of memristors mapping algebraic matrices, such as a neural network layer, multiple techniques have been proposed for overcoming limited device precision.
-For example, with bit slicing [11] by using $M$ levels memristor, $N=2^{2M}$ bits can be represented by connecting two devices on the same row.
+In the case of a model larger than what can be mapped in a single X-TIME chip, we foresee the possibility of multiple CPs in multiple chips orchestrating a reduction similar to the adder tree in the chip network.
+X-TIME is an inference accelerator, but models in production might be retrained and updated, due to data and/or concept drift, up to every 100 days [39].
+While ReRAM endurance might be limited to 105-108 cycles depending on the technology [40] given the large retraining cycle, the lifetime of the accelerator is significantly larger than the lifetime of typical systems, making ReRAM endurance not a strong requirement.
+4.2 Analog CAM with improved precision
+Figure 6: (a) Circuit schematic and (b) block diagram of a AND-type 4-bit 6T-2M analog CAM [15]. (c) Block diagram of a 4-bit OR-type analog CAM. (d) Block diagram of an 8-bit analog CAM.
+As a result of experiments in section 3 we found out that typically 8 bits precision, corresponding to 256 bins for each feature, is required for matching floating point (FP) accuracy inference of tree-based ML models.
+While analog CAM circuits offer the possibility of mapping arbitrary ranges, RRAM device reliability poses a challenge in supporting more than 16 levels or 4 bits [17].
+In the case of crossbar arrays mapping algebraic matrices, such as a neural network layer, multiple techniques have been proposed for overcoming limited device precision.
+For example, with bit slicing [12] by using two $M$ bits RRAMs, $N=2^{2M}$ bits can be represented by connecting two devices on the same row (Fig.
+LABEL:supp-si:fig:slicing(a)).
 In crossbar arrays, it is also straightforward to arbitrarily adjust the input precision by applying binary pulses and performing SHIFT and ADD operations at the output [12].
@@
 $2^{N_{bits}}$) rather than the number of bits.
-Thus given memristors with $M=4$ bits, 16 analog CAM cells are needed for representing $N=8$ bits, leading to an exponential overhead.
-Here, we propose a novel solution to increase the precision with the same base analog CAM cell circuit, which requires 2 analog CAM cells and 2 clock cycles for performing 8-bit search operation with $M=4$ bits memristors.
-To demonstrate, let us first take the case where the lower bound threshold $T_{L}$ is needed to be mapped and operated with $N=8$ bit of precision but we only have $M=4$ bits of precision in our analog CAM and memristor devices.
+Thus given RRAMs with $M=4$ bits, 16 analog CAM cells are needed for representing $N=8$ bits, leading to an exponential overhead (Fig.
+LABEL:supp-si:fig:slicing(b)).
+Here, we propose a novel solution to increase the precision with the same base analog CAM cell circuit, which requires 2 analog CAM cells and 2 clock cycles for performing 8-bit search operation with $M=4$ bits RRAMs.
+To demonstrate, let us first take the case where the lower bound threshold $T_{L}$ is needed to be mapped and operated with $N=8$ bit of precision but we only have $M=4$ bits of precision in our analog CAM and RRAMs devices.
 The lower bound threshold can be sub-divided into its most significant bit representation $T_{LMSB}$ and least significant bit representation $T_{LLSB}$, each one made of $M$ bits and written as $T_{L}=2^{M}T_{LMSB}+T_{LLSB}=16T_{LMSB}+T_{LLSB}$.
@@
 $\begin{split}[(q_{MSB}\geq T_{LMSB}+1)\vee(q_{LSB}\geq T_{LLSB})]\\ \wedge(q_{MSB}\geq T_{LMSB})\\ \wedge[(q_{MSB}<T_{HMSB})\vee(q_{LSB}<T_{HLSB})]\\ \wedge(q_{MSB}<T_{HMSB}+1).\end{split}$ (4)
-[…]
-TABLE I: Input to apply for doubling precision with time shifting
-Input Cycle 1 Cycle 2
-$q_{HLSB}$ $q_{LSB}$ $GND$
-$q_{LLSB}$ $q_{LSB}$ $VDD$
-$q_{HMSB}$ $q_{MSB}$ $q_{MSB}-1$
-$q_{LMSB}$ $q_{MSB}-1$ $q_{MSB}$
-Table I shows the required inputs for each DL analog bus wire to compute equation (4) in a 2-step search operation.
-With MAL pre-charged high, the first cycle evaluates the first bracket of equation (4) for the lower and upper bound (the OR operation).
-Without pre-charging MAL again, the second cycle evaluates the second term of each bound by inputting always care (i.e.
-always mismatch) values to the LSB sub-cell.
-Therefore, an AND operation between the outputs of cycle 1 and cycle 2 is computed on MAL following cycle 2, similar to a 2-step search approach previously demonstrated for increasing bit density in TCAMs [30].
-To the best of our knowledge, this is an entirely new approach and the proposed circuit operation is the most compact way of computing equation (4) with an analog CAM, and thus for doubling the bit precision.
-While other circuit designs may implement the same function in one clock cycle, these would have a significantly larger area.
-For example, one could separate cells performing the OR operation by connecting the discharge transistors in series rather than in parallel.
-This approach would also allow for selecting the precision of each feature to be either 4 or 8-bit, depending on the required precision.
-However, to perform such an operation in one cycle, three analog CAM cells with complex routing are needed.
-Given that analog CAM search latency is forecast to 100 ps [14], we found that performing two searches rather than one would not unduly increase the overall latency.
-Thus, we adopt this circuit formulation and operation which doubles the bit precision while doubling the area and analog CAM search latency.
-III-C Core pipeline
-In our ML experiments and hyperparameters search (see IV-A for details) we found out that $N_{leaves}=256$ per tree is typically sufficient to achieve high accuracy.
-Given that previous results have shown that analog CAM arrays with up to 128 rows do not encounter significant technology issues at 16 nm[14], each X-TIME core contains 2 stacked arrays, resulting in $2\times 128=256$ addressable words ($N_{leaves}$).
-We perform a survey on the size of typical tabular datasets [1], finding that usually $N_{feat}<100$ and often feature pre-processing reduces the number of $N_{feat}$ significantly before applying inputs to the decision tree ensemble.
-Previous results have shown that arrays with up to 100 columns have a very small access latency ($<$ 1ns) [14].
-Nevertheless, we include an outlier dataset with $N_{feat}=130$[31] to show how to handle larger input vector sizes with input vector segmentation, as will be described below.
-Each core contains 2 arrays of 65 columns, resulting in a width of $2\times 65=130$.
-Overall, each core has a total addressable analog CAM of macro array size $2\times 128\times 2\times 65=256\times 65$.
-The search operation is performed in series by groups of 65 features, while all the branches mapped to each queued array are computed in parallel.
-Multiple trees can be mapped to a single core, by arranging them in different portions of the analog CAM macro on the vertical axis.
-To avoid structural hazards, different considerations are made before applying new inputs to the core, as will be discussed below.
-When a feature vector is searched, a set of components with indices ranging from 00 to $64$ is sent to the first analog CAM logical array (aCAM1) while a second set of components with indices from $65$ to $129$ is sent to the second array (aCAM2).
-Comparison sequences on each tree branch are reconstructed by the AND operations made by aCAM1 and aCAM2 by selectively pre-charging only aCAM2 rows matched in aCAM1.
-Each analog CAM array has a latency of $\lambda_{CAM}=4$ cycles for performing the search operation.
-One cycle is used for pre-charging the MAL, two cycles for the search operation (MSB and LSB), and one cycle for latching the MAL sense amp, which leads to a total core latency of $\lambda_{C}=12$ cycles, with all other components (Buffer, MMR, SRAM, ACC) having a single cycle latency.
-In the case of up to four trees mapped to the core when an array finishes the search operation, it starts processing a new sample and no additional bubbles are required.
-The maximum ideal core throughput $\tau_{C}$, considering a clock frequency of 1GHz [14] and a relatively large number of samples $N_{s}$, can thus be computed by
-$\tau_{C}=\dfrac{N_{s}}{\lambda_{C}+\lambda_{CAM}(N_{s}-1)}\sim 250MSamples/s.$ (5)
-If more than 4 trees are mapped in the core, extra pipeline bubbles $N_{B}=N_{trees,core}$ should be inserted.
-In fact, the MMR needs to perform more than 4 iterations in order to generate the one-hot vector for accessing the leaves, stored in the SRAM.
-The core pipeline in this case reaches an ideal throughput of
-$\tau_{C}=\dfrac{N_{s}}{\lambda_{C}+N_{B}(N_{s}-1)}\sim 200MSamples/s\$ (6)
-To maintain such a high throughput, cores need to be fed with input data fast enough.
-On the other hand, the architecture should be able to adapt to multiple models (such as RF, XGBoost, and others, either classification or regression problems) and exploit the vast amount of cores for parallel computations.
-Thus we designed a programmable NoC able to handle multiple workloads.
-III-D Flexible connectivity
-We identified four possible examples of inference operation that cover all ensemble models and tasks previously described, and exploit all the possibilities of our architecture: (1) regression and binary classification, (2) multi-class classification, (3) regression and binary classification with input batching and (4) multi-class classification with input batching.
-In the case of regression and binary classification, leaves contain normalized partial predictions of each tree branch called logits, that are summed over the ensemble.
-In each core, the SRAM is programmed with logits that are summed locally in the core and in each network node inside the router.
-The final prediction is then sent to the CP which compares it with a threshold, typically 0 or 0.5, and sends the final result to the output.
-In the case of multiclass classification, logits should be summed class-wise and $N_{class}$ logits should be sent to the CP which performs an argmax computation to select the predicted class.
-Figure 6: Example of network configurations for regression/binary classification (a), multiclass classification (b) and regression/binary classification with input batching (c).
-(d) Schematic representation of X-TIME programming flow
-We propose a flexible/re-configurable NoC design (Fig.
-6(a-c)) to aggregate results based on model type.
-The inset of Fig.
-6(a) shows the downstream circuit schematic for the custom router.
-A buffer stores incoming leaves and has the output connected to a de-multiplexer (DEMUX) controlled by a configuration register.
-The DEMUX sends each leaf either to the accumulator circuit or directly to the output, with only logits belonging to the same class accumulated.
-Three examples of network configuration are shown in Fig.
-6 for supporting (a) regression or binary classification, (b) multi-class classification, and (c) regression or binary classification with input batching.
-In the first case, the X-TIME compiler maps $N_{trees}=8$ trees with a depth $D=7$ each, with $N_{trees,core}=2$ each core.
-The same inputs (blue solid and light blue dashed lines) are applied to all the cores in parallel and all the logits (solid purple lines) are summed together, being a regression/binary classification problem.
-The X-TIME compiler sets the configuration bit of all routers to ’1’ such that logits are always accumulated in the router before propagating.
-For multi-class classification (Fig.
-6(b)), the X-TIME compiler maps $N_{estimators}=2$ estimators each made of $N_{trees}=4$ (one for each class) with a depth $D=7$.
-Each core holds two trees belonging to the same class and the same inputs (blue solid and light blue dashed lines) are applied to all cores in parallel.
-The X-TIME compiler sets router configuration bits to 0, as it is not possible to perform an accumulation/reduction at the router level and the reduction must be done globally by the CP.
-Thus, logits for each class are transmitted independently (solid purple lines, dashed purple lines, solid pink lines, and dashed pink lines).
-Finally, for regression/binary classification with input batching (Fig.
-6(c)), the X-TIME compiler maps an ensemble of $N_{trees}=4$ with depth $d=7$ and to increase the throughput, replicates trees in multiple cores, such that cores C1 and C2 map the same trees of cores C3 and C4.
-Different inputs are applied to different cores, or groups of cores by appropriately addressing each feature vector.
-Core C1 and C2 receive a first input (solid blue line) while cores C3 and C4 receive a second input (dashed light blue).
-Logits can be accumulated at the router level only between trees belonging to the same input batch, as in the lower level of the architecture in Fig.
-6(c).
-The X-TIME compiler programs the lower level router configuration bits to ’1’ to enable local accumulation, and the higher level configuration bits to ’0’ to enable logits direct propagation.
-Input batching addresses one of the main issues of specialized accelerators in keeping resource utilization high [32].
-With a similar procedure, multiple unique models can be mapped to the accelerator, by assigning a different batch to each model.
-IV Evaluation Methodology
-IV-A Datasets and model training
-TABLE II: Datasets and models characterization
-Dataset ID Task Samples $\boldsymbol{N_{feat}}$ $\boldsymbol{N_{classes}}$ Model $\boldsymbol{N_{trees}}$ $\boldsymbol{N_{leaves,max}}$
-Churn modelling[33] 1 Binary Classification 10000 10 2 CatBoost 404 256
-Eye movements[34] 2 Multiclass Classification 10936 26 3 XGBoost 2352 256
-Forest cover type[35] 3 Multiclass Classification 581012 54 7 XGBoost 1351 231
-Gas concentration[31] 4 Multiclass Classification 13910 129 6 Random Forest 1356 217
-Gesture phase segmentation[36] 5 Multiclass Classification 9873 32 5 XGBoost 1895 256
-Telco customer churn[37] 6 Binary Classification 7032 19 2 XGBoost 159 4
-Rossmann stores sales[38] 7 Regression 610253 29 N/A XGBoost 2017 256
-For our experiments, we trained state-of-the-art models with three ML algorithms on a set of benchmark datasets (details in Table II).
-The algorithms included random forests (RF) from scikit-learn [28], and gradient boosted trees from XGBoost [17], and CatBoost [18].
+Analog CAMs cells arranged on the same MAL perform a bit-wise AND ($\wedge$) operation between them, thus a circuit addition is needed for performing the OR ($\vee$) operation.
+We start from the conventional analog CAM cell shown in Fig.
+6 (a).
+The lower bound (yellow, T1, $M_{l}o$) is encoded in a 1-transistor-1-resistor structure.
+If the input is high enough, it turns on T1, lowering the gate node of the pulldown T2 which stays off, resulting in a match.
+If the input voltage is low or alternative $M_{l}o$ conductance is high, T2 is turned on.
+Thus the lower bound checks if $SL_{1}>f(M_{l}o)$.
+A similar discussion can be applied to the higher bound (orange, T3, $M_{h}i$, T4, T5), in which an inverter (T4,T5) is added between the input voltage divider (T3,$M_{h}i$) and the pulldown (T6) to represent the higher bound and checking if $SL_{2}\leq(M_{h}i)$.
 Fig.
-6(d) shows a conceptual schematic of the workflow, whose output is a series of threshold maps (one for each tree) with the lower/upper bound to program in analog CAMs, the corresponding core/class address and leaf value for each node.
-IV-B Cycle-detailed simulator
-We developed a functional and cycle-detailed simulator [^1] https://github.com/HewlettPackard/X-TIME to evaluate latency, throughput, energy efficiency, area, and accuracy of the in-memory engine for tree-based model inference.
-The architecture simulator was developed using the Structural Simulation Toolkit (SST) [39] and estimates the performance on the benchmark datasets detailed above.
-SST, which is a parallel discrete-event simulator framework, can integrate modular components in a unified interface and simulate large-scale systems with MPI.
-It also provides interfaces for simulation components and components that perform actual functions such as processors, memory, and networks.
-For our system, the user-defined components emulate the cycle-level function of each block described in Fig.
-4, and data flow through the links connecting the components with specified delay.
-The cycle-level architecture simulation consisting of an H-tree NoC with 4096 cores and 1365 routers verified the error-free functional operation and the pipelined architecture without hazards, while also yielding system performance estimates.
-Figure 7: Area (a) and peak power (b) breakdown
-Fig.
-7 summarizes the power and area for the proposed architecture.
-Previous work generated layouts for the analog CAM at 16 nm  [14, 15], which included the analog CAM array, DAC, SA, and P-Ch area, and extracted power consumption using single cell parameters [15].
-The analog CAM model was previously validated by experimental measurements [14].
-Here, we simulate model mapping using memristor conductances from $1\mu S$ to $100\mu S$.
-The analog CAM model also includes parasitics that limit the CAM pre-charge and discharge time and the DAC operational frequency [14].
-Power and area values for register and logic are based on TSMC 16nm PDK.
-The MMR design is based on the matching token MMR topology [40] whose output is fed back to MAL registers enabling multiple iterations when multiple trees are mapped to the same core.
-Notably, the analog CAM arrays dominate area and power consumption, with peripheral components contributing negligibly.
-Also, the overall peak power consumption of 19 W is comparable to GPU idle power ($\sim$25 W).
-IV-C Profiling Tree Model Inference on NVIDIA GPUs, FPGA and SRAM based accelerators
-Figure 8: Schematic of the SRAM-based accelerator core used as digital baseline.
-To demonstrate the analog CAM performance capabilities, we compare our results to three alternative accelerator approaches: GPU, FPGA, and digital SRAM.
-As our first performance baseline, we profiled the inference step of the tree models and corresponding datasets shown in Table II on an NVIDIA V100 GPU.
-We measured each GPU kernel execution time with 5 repetitions to process data batches of increasing size, up to a saturation point, where throughput improvements were no longer observed.
-To estimate GPU upper bound performance, we isolated the kernel execution time of each model using NVIDIA’s nvprof tool, and did not include the memory transfer times between host (CPU) to device (GPU) or internal GPU transfers in our measured times.
-When possible, we profiled models using CUDA kernels from the NVIDIA RAPIDS Forest Inference Library (FIL).
-Moreover, we compared X-TIME with an FPGA baseline FAXID [8].
-FAXID is a custom hardware accelerator for low-latency ransomware detection using high-level synthesis (HLS) using binary classification models.
-As FAXID only supports binary classification, for our comparison we used a 1-vs-all approach to benchmark latency and throughput where we select one of the classes as class 0 and set the remaining classes to class 1.
-In FAXID the maximum depth was restricted to 6, as their ransomware detection accuracy did not increase with deeper trees.
-To keep the baseline results faithful to the FAXID HLS design, we optimized the hyperparameters of Table II for smaller depth and larger number of trees on FAXID to match accuracy with the other approaches.
-Finally, we compare X-TIME to a custom digital accelerator baseline using SRAM to isolate the performance benefits from the analog CAM specifically.
-This SRAM baseline uses a distributed architecture with each tree mapped to a small SRAM accessed serially by a floating point unit [5], as shown in Fig.
-8.
-Each SRAM word stores the node value and feature, the left and right children addresses, or the corresponding logit or class in the case of a leaf node.
-We considered the X-TIME network architecture and changed just the core operation (analog CAM or digital SRAM) to simulate the effect on throughput, of multiple accesses required for each tree.
-One might consider a digital TCAM to be another meaningful comparison.
-However, as previously demonstrated [41], using a traditional digital TCAM for performing on-access comparison would result in large overheads.
-In detail, $2^{8}$ TCAM columns would be needed to represent each feature with 8-bit precision, and thus tree inference of 129 features requires 33024 TCAM columns.
-Considering an 80-column TCAM macro has an access time of $\sim 500ps$ [42], inference with a TCAM would take $500\cdot 33024/80ps\approx 20us$ per tree, significantly larger than the few cycles analog CAM latency.
-V Results
-V-A Model inference performance
-Figure 9: (a) Accuracy for different training constraints and (b) relative accuracy as a function of memristor and DAC defects.
-In designing our HW architecture, we investigated the required bit accuracy and model types that the architecture should support to maintain state-of-the-art model accuracy.
-To assess these parameters, we compared model accuracy for different training parameters and model types to guide the X-TIME requirements (Fig.
-9(a)).
-First, we set our baseline, achieving state-of-the-art accuracy [43] by training all the models without constraints in the hyperparameters search (’Unconstrained’ in Fig.
-9(a)).
-We then optimized models to fit the hardware using the proposed 8-bit precision ACAM (‘X-TIME 8bit’: maximum $N_{trees}=4096$, $N_{leaves,max}=256,N_{bit}=8$), which resulted in limited accuracy loss compared to the unconstrained baseline case.
-To assess the need for 8-bit vs.
-4-bit analog CAMs, we optimized the model to fit hardware using a 4-bit ACAM [15], allowing the number of leaves to double, such that the area would be equivalent to ’X-time 8bit’(‘X-TIME 4bit’: maximum $N_{trees}=4096$, $N_{leaves,max}=512,N_{bit}=4$).
-Another option would double the maximum number of allowed trees, but we generally found that 4096 trees are high enough for matching unconstrained baseline accuracy.
-Our results showed a reduction in performance of $20\%$ in the case of regression on Rossmann store sales dataset [38] and $18\%$ in the case of multi-class classification on gas concentration dataset [31].
-Finally, we limited the optimizer to RF models only, which resulted in significant accuracy degradation for several datasets (label = ‘Only RF’).
-Given that it is not possible to train directly with 4-bit precision, the after-training quantization significantly decreased accuracy, demonstrating the need for supporting a broader range of model types beyond those demonstrated in previous works [44, 15].
-To assess the promise of this analog hardware approach in practical use, we also studied the impact of variations and defects of the analog hardware on model inference accuracy.
-The two most impactful sources of error to classification/regression results in this analog hardware are the known memristor conductance variations and DAC input variations to the aCAM.
-We considered experimental data from TaOx memristor and 16 nm analog circuits[15].
-Fig.
-9(b) shows the mean relative accuracy, defined as the ideal accuracy of Fig.
-9(a) over the defect-compromised accuracy as a function of the percentage of defects.
-We defined a defect as a 1-level random flip in the memristor conductance or DAC output voltage.
-The plot shows the average results across all classification datasets.
-We observe that these models and our architecture are fairly resilient to memristor variations, as individual errors do not have a large impact due to the ensemble nature of the model itself.
-With the aforementioned conductance window and a conservative standard deviation of $1\mu S$ it is expected to have a flip probability of $\sim 0.2\%$, corresponding to accuracy drop $<0.5\%$.
-On the other hand, while DAC variations also impact the accuracy, DACs are usually designed for operating up to 16 bits [45], thus realizing a stable/error-free 4-bit design is not particularly challenging.
-V-B Comparison with state-of-the-art
-Figure 10: Comparison of (a) latency and (b) throughput of this work, GPU, FPGA, and SRAM based accelerator from Fig. 8 (a) Throughput as a function of $N_{trees}$ and $D$.
-We benchmarked our proposed architecture on a wide range of models (corresponding to X-TIME 8bit of Fig.
-9) and dataset types (Table II), and compared latency and throughput (see Fig.
-10(a,b)) to GPU, FPGA and a digital accelerator based on small SRAMs and similar to Booster [5].
-When possible, we applied input batching and replicated trees to increase throughput.
+6 (b) shows a more compact circuit schematic for the analog CAM in Fig.
+6(a), with comparators driving the pulldown representing the 1T-1M and 1T-1M-1INV structures.
+We refer to this circuit as AND-type, given the AND operation between the lower and upper bound.
+Similarly, an OR-type circuit can be realized (Fig.
+6 (c)) [30], effectively performing an out of range matching.
+Finally, by combining the two structures, it is possible to realize the general circuit of Fig.
+6 (d) which can implement Eq.
+4, by applying opportunely inputs to the comparator (see Supplementary Information LABEL:supp-si:section:8bit-acam) Note that separate inputs for each cell are required for the programming operation, thus no extra wiring to the cell is required.
+5 Benchmark Results
+5.1 Comparison with state-of-the-art
+Figure 7: Comparison of (a) latency and (b) throughput of this work, GPU, FPGA, and fully digital ASIC accelerator from Fig. LABEL:supp-si:fig:digital_accelerator.
+We benchmarked X-TIME using our open-sourced cycle-approximate simulator [^1] https://github.com/HewlettPackard/X-TIME on a wide range of models and dataset types (Table 1), and compared latency and throughput (see Fig.
+7(a,b)) to GPU, specifically NVIDIA V100, FPGA [9] and a digital accelerator based on small SRAMs and similar to Booster [6] (Fig.
+LABEL:supp-si:fig:digital_accelerator).
+More details on the methodology can be found in Supplementary Information LABEL:supp-si:section:soa, LABEL:supp-si:section:methods and Table LABEL:supp-si:tab:benchmark.
+When possible, we applied input batching and replicated trees to increase throughput (Fig.
+LABEL:supp-si:fig:batching).
 As shown, our architecture produces drastically reduced latency, frequently to $\sim$100ns compared to GPU at 10 $\mu$s to ms, while also demonstrating significantly improved throughput by 10-120$\times$ across most datasets.
@@
 The SRAM-based accelerator has a linear dependence on $D$ (since multiple memory accesses are required) and $N_{trees}$ when more than a tree needs to be mapped in a core.
+Figure 8: Throughput as a function of number of trees $N_{trees}$ and depth $D$ for various accelerators.
+Note that in the case of X-TIME and the FPGA implementation, there is no dependence on the depth, removing the load imbalance issue, while only X-TIME shows no dependence on the number of trees.
 As an example, we studied the throughput as a function of $N_{trees}$ and $D$ for GPU, FPGA, SRAM-based accelerator, and X-TIME as shown in Fig.
-10(c), for a multiclass classification dataset [35], demonstrating the predicted trend.
+8, for a multiclass classification dataset [34], demonstrating the predicted trend.
 We note that the SRAM-based accelerator still suffers from load imbalance between trees, similar to GPU, in which trees with different depths have different inference times, which results in the need for synchronization before the reduction operation.
-The high throughput of X-TIME can produce an extremely low energy per decision, down to 0.3 nJ/Dec, highlighting the potential for energy-efficient applications.
-VI Conclusions
+As an energy comparison with GPU, we considered a worst-case scenario.
+Fig.
+LABEL:supp-si:fig:energy shows the energy saving in comparison by considering the peak power consumption (19 W) and idle power consumption (25 W) in the case of X-TIME and GPU, respectively.
+Even by assuming a full chip utilization for X-TIME and no energy consumption for the tree inference on GPU, X-TIME is up to 157$\times$ more energy efficient.
+5.2 Sensitivity to noise
+Figure 9: Accuracy and RMSE for different noise configurations
+Analog hardware is subject to noise and non-idealities.
+However, tree-based ML ensembles are resilient to variation thanks to the averaging mechanism happening between trees of multiple boosting rounds.
+Fig.
+9 shows the average accuracy of the models for different noise profiles, namely no noise (ideal), only ReRAM or DAC noise, and both noise sources, over 100 inference experiments.
+A printout of the data can also be found in Table  LABEL:supp-si:tab:noise.
+We characterized the noise of TaOx-based ReRAM devices at various conductance values ranging from $<0.1$ nS to 500 $\mu$S [27], by reading them 1000 times and extracting mean and standard deviation.
+We measured a maximum relative standard deviation of noise of $\sigma_{G}/G=0.1$.
+While DAC designs can reach higher precision, only 4-bits are needed for X-TIME inference.
+Nevertheless, we considered a pessimistic standard deviation of noise for the DAC $\sigma_{DAC}=50mV$.
+Results indicate no perturbation in the accuracy and RMSE when either ReRAM or DAC noise is applied, thanks to our hardware-aware training and inherent averaging mechanisms of ensemble models.
+In the case of regression and with both noise sources applied there is an RMSE perturbation, but still reduced compared to the 4-bit operation (Fig.
+4).
+Intuitively, with the thresholding mechanism of Analog CAM, it is possible to program conductances with a separation such that the probability of exceeding the threshold is relatively low, leading to good noise resistance performance.
+6 Conclusion
 In this work, we presented a novel architecture for accelerated tree-based ML inference with in-memory computing based on a novel increased precision analog non-volatile CAMs.
-Our approach maps decision trees to analog CAM ranges, which compute directly the matched results in parallel given an input feature vector.
+Analog CAMs enable novel applications of in-memory computing beyond matrix multiplies.
+Moreover, the required peripherals are significantly reduced compared to crossbar-based computing primitives.
+Our approach maps decision trees to analog CAM ranges, which directly compute the matched results in parallel, given an input feature vector.
 Leaf values are stored in a local SRAM and accumulated at the core level when multiple trees are mapped in the same core.
@@
 The on-chip network is re-configurable to map different models (regression, multi-class classification, etc) and inference techniques, such as input batching.
+An HW-aware training routine ensures to gets state of the art accuracy by leveraging quantization and optimizing the number of trees and maximum depth.
 Our results show significant improvement in latency – by almost four orders of magnitude, compared with an NVIDIA V100 GPU, at a throughput up to 120$\sim$ 120$\times$ 120 × higher across a wide range of dataset types and model topologies.
+Further, the worst case scenario showed up to 150×150$\times$150 × improvement in energy efficiency compared to GPUs with our approach.